### v1

# On Continuous Optimization for Constraint Satisfaction Problems

Abstract
Constraint satisfaction problems (CSPs) are fundamental in mathematics, physics, and theoretical computer science. While conflict-driven clause learning Boolean Satisfiability (SAT) solvers have achieved remarkable success and become the mainstream approach for Boolean satisfiability, recent advances show that modern continuous local search (CLS) solvers can achieve highly competitive results on certain classes of SAT problems. Motivated by these advances, we extend the CLS framework from Boolean SAT to general CSP with finite-domain variables and expressive constraints. We present FourierCSP, a continuous optimization framework that generalizes the Walsh-Fourier transform to CSP, allowing for transforming versatile constraints to compact multilinear polynomials, thereby avoiding the need for auxiliary variables and memory-intensive encodings. Our approach leverages efficient evaluation and differentiation of the objective via circuit-output probability and employs a projected gradient optimization method with theoretical guarantees. Empirical results on benchmark suites demonstrate that FourierCSP is scalable and competitive, significantly broadening the class of problems that can be efficiently solved by CLS techniques.
Introduction
Constraint Satisfaction Problems (CSPs) are foundational in mathematics, physics, and theoretical computer science. Given a set of variables with domains taken from finite sets, and a set of constraints on the variables, i.e., a CSP instance, solving the instance is to find an assignment of variables that satisfies all the constraints. Applications of CSP solving are in artificial intelligence (Baluta et al. 2019), database query (Kolaitis and Vardi 1998), operational research (Brailsford, Potts, and Smith 1999), statistical physics (Nagy et al. 2024), information theory (Golia, Juba, and Meel 2022), quantum computing (Vardi and Zhang 2022), and so on.
Most CSP solvers can be categorized into two types: systematic search and local search. Systematic methods are often built upon the Branch-and-Bound (BnB) framework, which systematically explores the solution space while pruning infeasible branches. Notably, SAT solving, driven by Conflict-Driven Clause Learning (CDCL), i.e., BnB with clause learning (Davis and Putnam 1960; Davis, Logemann, and Loveland 1962; Marques-Silva and Sakallah 1999), has shown remarkable engineering success, routinely handling industrial-scale problems involving millions of variables (Vardi 2014). Nevertheless, SAT solvers primarily handle Conjunctive Normal Form (CNF) formulas, necessitating Booleanizations (Huang 2008) and subsequent CNF encodings (Prestwich 2021), potentially leading to substantial increases in formula size and inefficiency.
Integer Linear Programming (ILP) provides a more general formulation than CNF clauses, permitting more compact CSP representations through linearization, and utilizes linear programming relaxations to prune the search space efficiently. However, ILP solvers generally exhibit weaker Boolean reasoning capabilities than SAT solvers (Elffers and Nordström 2018). Constraint programming (CP) offers a native approach, directly accepting constraints expressed via a broad range of predicates over diverse variable types (Krupke et al. 2024). The research community of CP has developed multiple open-source solvers (Schulte, Lagerkvist, and Tack 2006; Prud’homme and Fages 2022; Hebrard 2008; Kuchcinski and Szymanek 2013). However, they typically do not scale as efficiently as ILP solvers and are less effective at constrained optimization problems. The OR-Tools CP-SAT solver, a portfolio solver with SAT, CP, has effectively addressed many limitations of traditional CP solvers and emerged as a robust alternative to ILP solvers, demonstrating competitive performance across a wide range of problems (Perron, Didier, and Gay 2023).
Systematic search solvers are complete; they can prove unsatisfiability or optimality if provided sufficient computation time. In contrast, local search methods aim to discover feasible solutions efficiently without proving unsatisfiability. Local search methods can be categorized into discrete local search (DLS) and continuous local search (CLS). DLS iteratively changes the value of variables to maximize the number of satisfied constraints (Mitchell, Selman, and Leveque 1992; Selman et al. 1993). CLS, is initiated by UniSAT (Gu 1994) and followed by FourierSAT (Kyrillidis et al. 2020), GradSAT (Kyrillidis, Vardi, and Zhang 2021), and FastFourierSAT (Cen, Zhang, and Fong 2025). CLS-based SAT solvers have been recently actively studied as a novel line of SAT frameworks. These solvers transform SAT into a polynomial optimization problem over the real domain and apply gradient-based optimizers. The latest work on this line, named FastFourierSAT, reveals that CLS solvers can be highly parallelizable and amenable to GPU accelerators. Recent theoretical and engineering efforts help CLS solvers achieve state-of-the-art (SOTA) performance on certain benchmarks in SAT solving, e.g., parity learning with error problems (Crawford, Kearns, and Schapire 1994).
Contributions
Despite promising advancements, current CLS methods are restricted to Boolean domains, limiting applicability to problems with inherently finite, multi-valued variables. In this work, we extend the FourierSAT framework to general CSPs, overcoming the significant challenge of directly mapping finite-domain variable domains to continuous optimization domains. We introduce an orthonormal product Fourier basis, allowing each constraint to be represented as a linear combination of these bases. The constraints are then reduced to a continuous polynomial with randomized rounding, where the input represents the probability of an assignment, and the output corresponds to the expectation of the constraint being satisfied. The polynomial representation of a constraint, i.e., the Walsh-Fourier expansion (WFE), can have an exponential number of terms, and a naive implementation is memory- and time-inefficient. In this work, we show that the WFE can be represented by the circuit-output probability (COP) of a decision diagram, which can be efficiently evaluated and differentiated.
We implemented our approach as FourierCSP with projected gradient ascent, The benchmark instances include both constraint satisfaction and optimization problems, including task scheduling and graph coloring with hashing query problems. Our approach is accelerated with GPU parallelization and compared with SOTA SAT-based, ILP-based, and CP-based complete solvers. Despite the incompleteness, our solver obtained encouraging results on CSP solving. To the best of our knowledge, this paper is the first algorithmic work to introduce continuous optimization to CSP. This opens a new research direction, enabling the integration of future continuous optimization techniques and advanced parallel hardware.
Preliminaries
Constraint Satisfaction Problem
Let $X=\{x_{1},...,x_{n}\}$ be a set of variables, and $D=\{\mathrm{dom}(x_{1}),...,\mathrm{dom}(x_{n})\}$ be a set of finite domains, and $C=\{c_{1},...,c_{m}\}$ be a set of constraints. Every constraint $c_{i}\in{}C$ is described by an atom set $A_{i}\subseteq{}X$. The (finite-domain) CSP is to find an assignment of $X$ over $D$ such that a finite-domain formula, i.e., the conjunction of all constraints, $F=\bigwedge_{c_{i}\in{}C}c_{i}$, is satisfiable.
Walsh-Fourier Transform
We define a function with finite domain by $f:D\to\{0,1\}$, where 0 and $1$ present false, true of the finite constraint.
Walsh-Fourier Transform is a method for transforming a function with a finite domain into a multilinear polynomial. Given the Fourier bases, any function defined on a finite domain has an equivalent Fourier representation.
Definition 1 (Fourier Bases).
Given a function on a finite domain $f:D\to\{0,1\}$, we define the Fourier bases $\phi_{\alpha}$ by
$\phi_{\alpha}(X)=\prod_{i=1}^{n}\phi_{\alpha_{i}}(x_{i}),$
where $\alpha_{i}\in\mathrm{dom}(x_{i})$, and $\phi_{\alpha_{i}}$ is an indicator basis, i.e., $\phi_{\alpha_{i}}(x_{i})$ equals to $1$ if $\alpha_{i}=x_{i}$, otherwise equals to 0.
Theorem 1 (Walsh-Fourier Expansion (O’Donnell 2014)).
Having fixed the Fourier bases, $f$ is uniquely expressible as
$f(X)=\sum_{\alpha\in{}D}\left(\hat{f}(\alpha)\cdot\phi_{\alpha}(X)\right)$
where $\hat{f}(\alpha)\in\mathbb{R}$ is called Fourier coefficient, and it satisfies
$\hat{f}(\alpha)=\mathop{\mathbb{E}}_{X\sim D}\left[f(X)\cdot\phi_{\alpha}(X)\right].$
Theoretical Framework
In this section, we first recap the previous Continuous Local Search (CLS) for Boolean SAT solving. We then demonstrate how our proposed solver, FourierCSP, can extend the CLS framework to general constraint satisfaction problems. Proofs are delayed to the supplemental material.
Continuous Local Search
CLS-based solvers define a continuous objective function, the optima of which encode the solutions to the original problem defined on discrete variables. The CLS framework is described in Algorithm 1
Algorithm 1 Continuous Local Search
Input: A formula $F$ with a constraint set $C$
Output: A discrete assignment $X\in D$
1: Sample $P$ from $\tilde{D}$
2:for $j=1,\cdots,\texttt{max\_iter}$ do
3:  Search for a local optimum $P^{*}$ of $\mathcal{F}(P)$
4:  Discretized assignment: $X=\mathcal{R}(P^{*})$
5:  if $\mathcal{F}(X)=|C|$ then
6:   return $X$
7:  end if
8:end for
9:return $X$ with the highest $\mathcal{F}(X)$
Definition 2 (Objective).
Given a formula $F=\bigwedge_{c\in{}C}c$, every constraint $c\in{}C$ can be transformed to a function by Theorem 1. We denote the corresponding expansion as $f_{c}$ and the objective function w.r.t. $F$ by
$\mathcal{F}=\sum_{c\in{}C}f_{c}.$
FourierSAT, FastFourierSAT, and GradSAT are three gradient-based CLS variants. They transform Boolean constraints to continuous functions, called WFE via the Walsh-Fourier transform (O’Donnell 2014). However, the WFE of $n$ variables has at most $2^{n}$ terms. FourierSAT shows that, when the Boolean constraints are symmetric, e.g., XOR, cardinality, not-all-equal constraints, they can be concisely represented by $n$ elementary symmetric polynomials (Kyrillidis et al. 2020). FastFourierSAT demonstrates that elementary symmetric polynomials can be computed by convolutions, which correspond to multiplications in the frequency domain (Cen, Zhang, and Fong 2025). When the constraint is not symmetric, e.g., Pseudo-Boolean constraint, GradSAT can concisely represent the WFE by the COP of a binary decision diagram (BDD) (Kyrillidis, Vardi, and Zhang 2021).
Limitations of Previous CLS Solvers
The entire lineage of CLS solvers, from the 1990s approaches (Gu 1994) to modern variants like FourierSAT, FastFourierSAT, and GradSAT, is exclusively designed for Boolean constraints. While they perform well on certain problem classes, these solvers often fail to scale effectively to large or complex CSP instances. This limitation becomes particularly acute for finite-domain problems, which require conversion to potentially larger Boolean formulas. This limitation fundamentally restricts their practical application for real-world problems that naturally require finite-domain representations.
FourierCSP
To address the above limitation, we propose FourierCSP, a CLS framework that directly supports finite-domain CSPs without requiring Booleanization. We start by applying randomized rounding to the generalized form of WFE in Theorem 1. With the randomized rounding, we can reduce a constraint satisfaction problem to a continuous optimization problem. Subsequently, we reveal that the objective function is equivalent to the circuit-output probability of a decision diagram. Hence, the computational complexity of objective evaluation, i.e., one of the key bottlenecks for continuous optimizers, scales with the size of the decision diagram. Moreover, we derive practical convergence bounds that gradient-based methods can achieve within our framework.
Reduction by Randomized Rounding
Theorem 1 illustrates how to expand discrete functions to multilinear polynomials through Walsh-Fourier expansion. The resulting Fourier coefficients $\hat{f}(\alpha)$ provide valuable insights into analyzing the discrete function properties. However, the expanded function $f$ remains discrete and non-differentiable given the Fourier Bases in Definition 1, limiting its utility for gradient-based optimization.
To enable continuous optimization, we introduce a randomized rounding $\mathcal{R}$ that relaxes discrete variable assignments into a continuous domain. Consider a set of finite domain variables $X=\{x_{1},...,x_{n}\}$. The vector probability space of $x_{i}$ is a simplex denoted by
$\tilde{D}_{i}:=\left\{p_{i,j}\Bigg|\min_{j\in\mathrm{dom}(x_{i})}p_{i,j}\geq{}0,\sum_{j\in\mathrm{dom}(x_{i})}p_{i,j}=1\right\}.$
The collection of all simplexes forms a continuous search space $\tilde{D}=\left(\tilde{D}_{1}\times\cdots\times\tilde{D}_{n}\right)$.
Definition 3 (Randomized Rounding).
Given a finite domain $D$. The randomized rounding function, denoted by $\mathcal{R}:\tilde{D}\to{}D$ is defined by
$\mathbb{P}[\mathcal{R}(P)_{i}=j]=p_{i,j}$
for all $j\in\mathrm{dom}(x_{i})$, $p_{i}\in\tilde{D}_{i}$ and $i\in[n]$.
Intuitively, if $p_{i,j}$ is closer to $1$, then $\mathcal{R}(a)_{i}$ is more likely to be rounded to $j$. Using this randomized interpretation, we can define a probability distribution over full assignments by rounding $\mathcal{R}$ of each variable independently.
Definition 4 (Probability Space).
We define a probability space on a discrete matrix w.r.t. to real point $P\in\tilde{D}$, denoted by $\mathcal{S}:D\to\tilde{D}$, as:
$\mathcal{S}_{P}(X)=\mathbb{P}[\mathcal{R}(P)=X]=\prod_{i=1}^{n}p_{i,x_{i}},$
for all $X\in{}D$, with respect to the randomized-rounding function $\mathcal{R}$. We use $X\sim\mathcal{S}_{P}$ to denote that $X$ is sampled from the probability space $\mathcal{S}_{P}$.
This probability space enables us to define a continuous extension of any discrete function by taking its expectation over a randomized assignment.
Theorem 2 (Expectation).
Given a Walsh-Fourier expansion $\mathcal{F}$, for a real point $P\in\tilde{D}$, we have
$\mathop{\mathbb{E}}_{X\sim\mathcal{S}_{P}}[\mathcal{F}(X)]=\sum_{c\in{}C}\sum_{\alpha}\left(\hat{f}_{c}(\alpha)\cdot\prod_{i=1}^{n}p_{i,\alpha_{i}}\right),$ (1)
where the expectation is a multilinear polynomial.
Hence, the satisfiability problem of a CSP can be reduced to a continuous optimization problem by Theorem 3.
Theorem 3 (Reduction).
Given a constraint satisfaction problem of formula $F$, which is the conjunction of all constraints $c\in{}C$. $F$ is satisfiable if and only if
$\max_{P\in\tilde{D}}\left(\mathop{\mathbb{E}}_{X\sim\mathcal{S}_{P}}[\mathcal{F}(X)]\right)=|C|.$ (2)
Circuit-Output Probability
Theorem 3 reveals that CSP solving can be reduced to a continuous optimization problem. The performance of gradient-based optimization crucially depends on the efficiency of evaluating and differentiating the objective function. However, the complexity of evaluation and differentiation is in #P-hard, i.e., Eq. (1) has at most exponentially many terms, in general. Motivated by (Thornton and Nair 1994), we reformulate Eq. (1) to circuit-output probability (COP), which admits more tractable computation when a structured representation can encode the underlying constraint, i.e., decision diagrams.
Definition 5 (Circuit-Output Probability).
Let $f_{c}$ be a discrete function over a set $X$ of finite domain variables. Let $P$ be the variable input probabilities. The circuit-output probability problem of $f_{c}$ w.r.t $P$, denoted by $\texttt{COP}_{c}(P)$, is the probability of $f_{c}$ outputs $1$ given the value of each variable independently sampled from a multimodal distribution, i.e.,
$\texttt{COP}_{c}(P)=\sum_{X\in D}f_{c}(X)\prod_{i=1}^{n}p_{i,x_{i}}.$
Corollary 1.
Given a constraint $c$ defined on a finite domain $D$, for a real point $P\in\tilde{D}$, we have
$\mathop{\mathbb{E}}_{X\sim\mathcal{S}_{P}}[f_{c}(X)]=\texttt{COP}_{c}(P).$
Let $\mathcal{C}(P)=\sum_{c\in{}C}\texttt{COP}_{c}(P)$, we have
$\mathcal{C}(P)=\mathop{\mathbb{E}}_{X\sim\mathcal{S}_{P}}[\mathcal{F}(X)]$
Corollary 1 implies that it is sufficient to evaluate Eq. (1) by computing the COP of each constraint. However, the complexity of calculating the COP is still in #P-hard for general representations. Nevertheless, if a constraint is represented by a decision diagram $G(V,E)$, then the COP on it can be obtained in $\mathcal{O}(|E|)$, where $|E|$ is the number of edges in $G$. Prior work demonstrates that BDD can encode many Boolean constraints in compact size (Kyrillidis, Vardi, and Zhang 2021; Sasao, Fujita et al. 1996; Aavani 2011).
The multiple-valued extension of BDD is called multi-valued decision diagram (MDD) (Bryant 1995). The MDD can be versatile in representing many constraints through the translation from table constraints (Cheng and Yap 2010). Similar to BDD, two MDDs can be combined through an APPLY operation (Bryant 1986; Perez and Régin 2015), offering flexibility in encoding versatile constraints. The MDDs can be reduced by merging the nodes having the same set of outgoing neighbors associated with the same label (Cheng and Yap 2010; Perez and Régin 2015). Moreover, the size can be further reduced by sifting (Miller and Drechsler 2003). A compact-size decision diagram means that the COP on it can be evaluated and differentiated efficiently.
Lemma 1 (Top-Down Evaluation).
Let $G(V,E)$ be the MDD of constraint $c$, run Algorithm 2 on $G$ with a real assignment $P\in\tilde{D}$. Let $X\in D$ randomly rounded from $P$ by $\mathbb{P}[x_{i}=j]=p_{i,j}$ for all $j\in\mathrm{dom}(x_{i})$. For each $v\in V$, we have
$m_{td}[v]=\mathbb{P}_{M,a}[v],$
where $\mathbb{P}_{G,a}[v]$ is the probability that the node $v$ is on the path generated by $b$ on $G$. Especially,
$m_{td}[\texttt{true}]=\texttt{COP}_{c}(P_{a}),$
i.e., Algorithm 2 returns the COP of $G$ in time $O(|E|)$.
Algorithm 2 Top-Down Traversal of MDD
Input An MDD $G(V,E)$ representation of constraint $c$, a real assignment $P\in\tilde{D}$.
Output Circuit-output probability $\texttt{COP}_{c}(P)$.
1: Let $m_{td}[v]$ be the top-down message of each node $v\in{}V$, where the root node is initialized as 1, otherwise 0.
2:for $v\in{}\texttt{sorted}(V)$ do
3:  for $u=\mathrm{child}(v)$ do
4:   $m_{td}[u]+=p_{\mathrm{id}(v),\mathrm{label}(v,u)}\cdot m_{td}[v]$
5:  end for
6:end for
7:return $m_{td}[true]$
Alternatively, traversing the MDD bottom-up can also compute the circuit-output probability.
Lemma 2 (Bottom-Up Evaluation).
Let $G(V,E)$ be the MDD of constraint $c$, run Algorithm 2 on $G$ with a real assignment $P\in\tilde{D}$. For each $v\in V$, let $f_{c,v}$ be the sub-function of the sub-MDD generated by regarding $v$ as the root. The following holds:
$m_{bu}[v]=\texttt{COP}_{c,v}(P),$
i.e., Algorithm 3 returns the COP of $G$ and in time $O(|E|)$.
Algorithm 3 Bottom-Up Traversal of MDD
Input An MDD $G(V,E)$ representation of constraint $c$, a real assignment $P\in\tilde{D}$.
Output Circuit-output probability $\texttt{COP}_{c}(P)$.
1: Let $m_{bu}[v]$ be the bottom-up message of each node $v\in{}V$, where the true terminal node is initialized as 1, otherwise 0.
2:for $v\in{}\texttt{sorted}(V,\texttt{reverse=True})$ do
3:  for $u\in\mathrm{child}(v)$ do
4:   $i=\mathrm{id}(v)$
5:   $m_{bu}[v]+=p_{\mathrm{id}(v),\mathrm{label}(v,u)}\cdot m_{bu}[u]$
6:  end for
7:end for
8:return $m_{bu}[root]$
The main idea of differentiation is from probabilistic inference (Pearl 1988; Shafer and Shenoy 1990), which traverses the graph model top-down and bottom-up. The following theorem shows that the partial derivative of COP can be obtained by top-down and bottom-up messages from Algorithm 2 and 3.
Theorem 4 (Derivatives).
After running Algorithm 2 and 3 on MDD for constraint $c$ with a real assignment $P\in\tilde{D}$, the partial derivative of COP on $p_{i,j}$ for all $j\in\mathrm{dom}(x_{i})$ can be computed by
$\frac{\partial\texttt{COP}_{c}(P)}{\partial p_{i,j}}=\sum_{v\in\{u:id(u)=i\}}m_{td}[v]\cdot m_{bu}[v.j]$ (3)
Projected Gradient
Since Eq. (1) is defined only over the constrained domain $\tilde{D}$, a single gradient step may result in an intermediate point that falls outside this domain. Therefore, each gradient step, i.e., Eq. (4), must be followed by a projection, i.e., Eq. (5), back onto $\tilde{D}$ to ensure feasibility.
$P_{t+1}^{\prime}$ $=P_{t}+\eta\cdot\nabla\mathcal{C}(P_{t})$ (4)
$P_{t+1}$ $=\mathrm{proj}_{\tilde{D}}(P_{t+1}^{\prime})$ (5)
Based on the randomized rounding in Definition 3, we define the projection for Eq. (5) as follows.
Definition 6 (Projection).
The Euclidean projection of $Q\in\mathbb{R}^{|\mathrm{dom}(x_{1})|\times\cdots\times|\mathrm{dom}(x_{n})|}$, onto $P\in\tilde{D}$, is defined as
[…]
where $I_{\tilde{D}}$ is an indication function. $I_{\tilde{D}}(P)$ is $\infty$ (resp. 0) when $P\notin\tilde{D}$ (resp. $\in$).
The projection operation $\mathrm{proj}_{\tilde{D}}(Q)$ can be computed efficiently by separately projecting each $q_{i}$ onto its respective simplex $\tilde{D}_{i}$, leveraging the product structure of $\tilde{D}$. Each sub-projection corresponds to a standard simplex projection problem (Chen and Ye 2011). By substituting Eq. (5) into Eq. (4), we have the gradient mapping:
$g(P)=\frac{1}{\eta}\left(\mathrm{proj}_{\tilde{D}}(P+\eta\cdot\nabla\mathcal{C}(P))-P\right).$
Lemma 3 and Theorem 5 demonstrate that projected gradient ascent is guaranteed to converge to a point where the gradient mapping is small, i.e., $g(P)||_{2}^{2}\leq\epsilon$.
Lemma 3.
Let $\mathcal{C}$ be a multilinear polynomial. Let $n$ be the number of variable, then $N=\sum_{i=1}^{n}|\mathrm{dom}(x_{i})|$. Let $g(\cdot)$ be the projected gradient mapping onto the simplices $\tilde{D}$. Given $y,z\in\tilde{D}$, if the step size satisfies $\eta\leq\frac{1}{|C|\sqrt{N}}$, then the projected gradient step $y=z+\eta{}g(z)$ satisfies:
$\mathcal{C}(y)-\mathcal{C}(z)\geq\frac{\eta}{2}\|g(z)\|^{2}$
Theorem 5 (Convergence Speed).
With the step size $\eta=\frac{1}{|C|\sqrt{N}}$, the projected ascent direction will converge to a $\epsilon$-projected-critical point in $\mathcal{O}\left(\frac{|C|^{2}\sqrt{N}}{\epsilon^{2}}\right)$ iterations.
The projected gradient moves the assignment toward a point associated with a higher function value, and eventually it will converge to a local maximum where the gradient mapping is small. If local maxima were to exist in the interior of the feasible region, i.e., in $\tilde{D}\setminus\delta\tilde{D}$, the quality of the corresponding decoded solution would not be guaranteed; however, we establish the non-existence of such interior local maxima in Lemma 4. Consequently, this result directly implies Theorem 6.
Lemma 4.
Let $\tilde{D}_{\perp}$ be a hyperplane with $\vec{1}$ as the normal vector. For every non-constant $\mathcal{C}$ and a point $P\in\tilde{D}\backslash{\delta\tilde{D}}$, there exist a region size $\epsilon>0$ and two directions, $q^{+},q^{-}\in\tilde{D}_{\perp}$ such that for all $\eta\in(0,\epsilon)$, the following holds.
$\mathcal{C}(P+\eta{}q^{+})>\mathcal{C}(P)$
Theorem 6 (Optimality).
Any interior point $P\in\tilde{D}\backslash{}\delta\tilde{D}$ cannot be a local maximum of $\mathcal{C}(P)$. The maximal points are only attained from the boundary $\delta\tilde{D}$.
To further illustrate these theoretical results, we include a selection of test cases in the supplementary materials, demonstrating how the projected gradient consistently converges to the boundary $\delta\tilde{D}$ across various scenarios.
Implementation Techniques
The primary computational workload of FourierCSP consists of repeatedly executing Algorithm 2 and Algorithm 3 for each constraint during optimization. Since the COP of individual constraints is independent of one another, these computations can be parallelized. Specifically, both algorithms can be compiled into kernel functions and executed concurrently, thereby accelerating the projected gradient step via hardware-level parallelism. Additionally, we leverage autodiff to compute gradients. As a result, it suffices to construct only the top-down evaluation trace. The top-down messages at the terminal nodes of the decision diagram are complementary, i.e., $m_{td}[\texttt{true}]+m_{td}[\texttt{false}]=1$. This allows us to compute the top-down message for only one terminal node while skipping all nodes that do not contribute to its computation. When the COP expression is sufficiently compact, it can be evaluated in closed form without traversing the decision diagram, thereby further reducing computational overhead.
Experimental Results
In this section, we compare our tool, FourierCSP, with systematic and local search solvers. The objective of our experimental evaluation is to answer three research questions:
RQ1. Given their Boolean-encoded formulas of CSP, how well does FourierCSP perform in solving them?
RQ2. How does FourierCSP compare to SAT, ILP, and CP solvers in solving constraint satisfaction problems that the constraints have structural patterns?
RQ3. How does FourierCSP perform relative to SAT, ILP, and CP solvers in solving constraint optimization problems that the constraints have structural patterns?
Benchmarks and Encodings
To fully exploit the potential of Fourier expansion-based CLS, benchmark instances should ideally be symmetric, i.e., their constraints should exhibit similar structural patterns. This symmetry enables efficient GPU acceleration by allowing the solver to fully utilize parallelism, achieving up to 100+ times speedup in gradient computation compared to CPU (Cen, Zhang, and Fong 2025). The benchmarks include constraint satisfaction and optimization problems. The first benchmark is task scheduling problems, which can be useful for the trace scheduler of multicore hardware systems. For each combination of $N_{\text{worker}}\in\{32,64,128,256,512\}$ and $m\in\{4,8,16\}$, we set the number of jobs to $N_{\text{job}}=m\cdot N_{\text{worker}}$, and generate 10 random instances accordingly. The second benchmark, graph coloring with hashing queries, is an optimization problem motivated by approximate model counting (Pote, Meel, and Yang 2025), where hash functions are employed to partition the solution space. Specifically, we use graph coloring as the underlying problem and add parity (modulo 2) constraints to the formula to serve as hash functions. For each $N_{\text{node}}\in\{512,1024,2048,4096\}$ and $N_{\text{color}}\in\{8,16,32,64\}$, we generate 10 random instances. Additionally, $N_{\text{node}}$ modulo 2 constraints are added to each instance.
Benchmark instances are linearized for ILP solvers and Booleanized for Boolean SAT solvers. Further details on the benchmarks and the encodings used for different solvers can be found in the appendix, where we show that the Booleanization will naturally lead to pseudo-Boolean and XOR constraints. In particular, further encoding the largest Booleanized instance into CNF (Ignatiev, Morgado, and Marques-Silva 2018; Li 2000) yields a formula comprising over 160 million variables and 710 million clauses. Although modern SAT solvers are highly optimized and capable of solving large benchmarks, such encodings are extremely memory-inefficient; for instance, the corresponding CNF text file is around 16 GB in size.
Solver Competitors
(1) FourierSAT (Kyrillidis et al. 2020): A CLS solver designed for Boolean formulas that are conjunctions of symmetric Boolean constraints. (2) FourierMaxSAT (Cen, Zhang, and Fong 2025): An extension of FourierSAT incorporating an SAT solver for warm-start initialization. (3) LinPB (Yang and Meel 2021): A SAT solver for linear pseudo-Boolean and XOR constraints. It employs RoundingSAT, which extends CDCL to pseudo-Boolean reasoning, as the underlying pseudo-Boolean solver. XOR constraints are supported via Gauss-Jordan elimination, adapted from CryptoMiniSAT. (4) Gurobi-ILP (Gurobi Optimization, LLC 2024): An ILP solver that uses branch-and-bound, cutting planes, and heuristics to handle integer linear constraints efficiently. The solver also leverages preprocessing and parallel computation to enhance performance. We use version 11.0.3. (5) OR-Tools CP-SAT (Perron, Didier, and Gay 2023): A hybrid solver that integrates a clause-learning SAT core, constraint programming, and linear relaxation within a unified model to solve combinatorial problems using advanced techniques; it won gold medals in all categories of the MiniZinc Challenge 2024. This solver won gold medals in all categories in the MiniZinc Challenge 2024. We use version 9.11. (6) ParaILP (Lin et al. 2024): A parallel discrete local search solver for ILP.
Experimental Setups
[…]
Answer to RQ1.
Since FourierCSP is the multi-valued extension of FourierSAT, the semantics of solving hybrid SAT formulas remain equivalent to those of FourierSAT. Prior research indicates that FourierSAT demonstrates strong performance on benchmarks natively formulated as hybrid SAT problems. However, our experiments reveal a significant limitation: FourierSAT fails to solve any instances encoded from both scheduling and graph coloring benchmarks. This observation underscores a critical gap in the capability of FourierSAT when handling problems not inherently structured as conjunctions of symmetric Boolean constraints. Consequently, it highlights the necessity of extending the solver to a more general domain. Specifically, the results motivate the development of FourierCSP, a multi-valued extension of FourierSAT, which is designed to address precisely these limitations by supporting more general, i.e.. multi-valued formulations.
Figure 1: Results on task scheduling problems. The numbers in the legend are the average PAR-2 score of the solvers.
Figure 2: Results on graph coloring with hashing query problems. The numbers in the legend are (the average relative score, #Win of the solvers).
Answer to RQ2.
We assess solver performance on constraint satisfaction problems using the PAR-2 (penalized average runtime 2) score, a metric balancing runtime efficiency and robustness by imposing penalties on timeouts.
$\text{PAR-2}=\frac{1}{N}\sum_{i=1}^{N}\left(t_{i}\cdot{}\mathbf{1}[t_{i}\leq T]+2T\cdot\mathbf{1}[t_{i}>T]\right)$
where $\mathbf{1}(\cdot)$ is the indicator function and $T$ is the time limit.
Figure 1 presents the results on the first benchmark, constraint satisfaction problems. Specifically, FourierCSP and ParaILP represent two parallel local search solvers based on continuous and discrete optimization methods, respectively. Notably, under a timeout limit of 1000 seconds, FourierCSP consistently solves all 150 instances, substantially surpassing ParaILP, which solves only 60. This significant performance gap can be attributed to the differences in the search methods. Discrete local search methods require that each assignment strictly increase the number of satisfied constraints, often leading to plateaus, particularly for constraints that are long or structurally complex, where small perturbations to discrete variable assignments may not immediately yield progress. In contrast, CLS methods can exploit fine-grained improvements: the optimizer advances as long as the continuous objective function increases, even marginally.
[…]
Answer to RQ3.
We evaluate this benchmark using two metrics: the relative score and the number of instances for which a solver attains the best objective value (#Win). The relative score is defined as follows.
$\text{score}(s,i)=\frac{1+c_{i}^{s}}{1+c_{i}^{*}}$
where $c_{i}^{s}$ denotes the result obtained for instance $i$ by algorithm $s$, and $c_{i}^{*}$ denotes the best-known result.
Notably, this benchmark includes a substantial number of modulo constraints, which are known to be challenging for CP-SAT (Krupke et al. 2024). While modulo constraints remain beyond the expressive or algorithmic reach of most conventional solvers, FourierCSP can natively accommodate them through its uniform Fourier expansion framework. As in Figure 2, FourierCSP outperforms all competing methods on this benchmark, achieving a relative score of 0.98 and attaining the best solution for 84 out of 160 instances In contrast to its competitive performance on the previous benchmark, CP-SAT’s relative score is 0.18 lower than that of Gurobi-ILP for these instances. This relative underperformance can likely be attributed to CP-SAT’s inefficiency in processing modulo constraints.
Under binary logarithm encoding, a modulo 2 constraint can be easily Booleanized to an XOR constraint. Although LinPB can natively solve XOR constraints using Gaussian elimination, it achieves a relative score of 0.16 on this benchmark. Nevertheless, LinPB contributes more to the virtual best solver than CP-SAT, as it attains the best solution for 17 instances. Notably, formulas with parity constraints are known to be challenging for DLS solvers (Crawford, Kearns, and Schapire 1994). It is unsurprising that ParaILP fails to provide valid solutions for any instance in this benchmark.
Previous work demonstrated that FourierSAT performs competitively on graph problems when the encoded formulas comprise exclusively symmetric Boolean constraints (Cen, Zhang, and Fong 2025). In contrast, the Booleanized formulas used in this benchmark consist primarily of pseudo-Boolean constraints, which would require further symmetrization into disjunctive clauses to fit the FourierSAT framework. This additional encoding step increases the complexity, making the problem significantly more challenging for FourierSAT, which does not provide valid solutions for any instance in this benchmark. While FourierSAT encounters difficulties when addressing formulas containing a large number of clauses, FourierMaxSAT adopts a hybrid approach. It leverages an SAT solver to solve part of the formula, i.e., the clauses, and subsequently uses CLS to improve the solution, i.e., by attempting to satisfy more XOR constraints. Eventually, FourierMaxSAT achieved a relative score of 0.94.
Moreover, we observe that the instances for which FourierCSP attains the best solutions are typically those with larger problem sizes. This observation suggests a qualitative distinction in scalability: while systematic search solvers tend to be effective for smaller instances, FourierCSP exhibits improved scalability and maintains solution quality as instance size increases.
Conclusion and Future Directions
In this paper, we introduced FourierCSP, a novel continuous optimization framework for constraint satisfaction problems. By extending the Walsh-Fourier expansion to finite-domain variables, our approach can handle versatile constraints without the need for memory-intensive encodings or auxiliary variables. The equivalence between circuit-output probability and Walsh-Fourier expansion allows us to evaluate and differentiate the objective function efficiently, enabling gradient-based optimization. We adopted a projected gradient approach for our optimization framework, which is supported by theoretical guarantees on convergence. Empirical evaluations on benchmark suites demonstrated that FourierCSP is scalable and performs competitively. The experimental results indicate that FourierCSP is a substantial extension of FourierSAT, broadening the class of problems that can be efficiently solved using CLS techniques.
The performance of Fourier transform-based CLS relies heavily on GPU parallelism. An important direction for future work is to investigate how this parallelism can be effectively leveraged when constraints exhibit non-structural patterns. This would further generalize FourierCSP, enabling its application to a broader range of CSP problems.
Acknowledgments
We thank Moshe Vardi for the helpful and insightful comments and discussions that contributed to this work. This work was supported in part by the National Research Foundation under the Prime Minister’s Office, Singapore, through the Competitive Research Programme (project ID NRF-CRP24-2020-0002 and NRF-CRP24-2020-0003); in part by the Ministry of Education, Singapore, through the Academic Research Fund Tier 2 (project ID MOE-T2EP50221-0008); and in part by the National University of Singapore, through the Microelectronics Seed Fund (FY2024). Zhiwei Zhang is supported in part by NSF grants (IIS-1527668, CCF1704883, IIS1830549), DoD MURI grant (N00014-20-1-2787), Andrew Ladd Graduate Fellowship of Rice Ken Kennedy Institute, and an award from the Maryland Procurement Office.
Appendix A Technical Proof
Proof of Theorem 2
Let $P\in\tilde{D}$ and $X\in{}D$.
$\mathop{\mathbb{E}}_{X\sim\mathcal{S}_{P}}[\mathcal{F}(X)]$ $=\mathop{\mathbb{E}}_{X\sim\mathcal{S}_{P}}[\sum_{c\in{}C}f_{c}(X)]$
$=\sum_{c\in{}C}\mathop{\mathbb{E}}_{X\sim\mathcal{S}_{P}}[f_{c}(X)]$
$=\sum_{\alpha\in{}D}\hat{\mathcal{F}}(\alpha)\cdot\mathop{\mathbb{E}}_{X\sim\mathcal{S}_{P}}[\phi_{\alpha}(X)]$ (7)
Since $\phi_{\alpha}(X)\in\{0,1\}$ By linearity of the expectation, we have
$\mathop{\mathbb{E}}_{X\sim\mathcal{S}_{P}}[\phi_{\alpha}(X)]$ $=1\cdot\mathop{\mathbb{P}}_{X\sim\mathcal{S}_{P}}[\phi_{\alpha}(X)=1]+0\cdot\mathop{\mathbb{P}}_{X\sim\mathcal{S}_{P}}[\phi_{\alpha}(X)=0]$
$=\mathop{\mathbb{P}}_{X\sim\mathcal{S}_{P}}[\phi_{\alpha}(X)=1]$
$=\mathop{\mathbb{P}}_{X\sim\mathcal{S}_{P}}\left[\prod_{i=1}^{n}\phi_{\alpha_{i}}(x_{i})=1\right].$ (8)
As $\phi_{\alpha_{i}}(\cdot)$’s are indication function,
$\mathop{\mathbb{P}}_{X\sim\mathcal{S}_{P}}\left[\prod_{i=1}^{n}\phi_{\alpha_{i}}(x_{i})=1\right]$ $=\prod_{i=1}^{n}\mathop{\mathbb{P}}_{X\sim\mathcal{S}_{P}}[\phi_{\alpha_{i}}(x_{i})=1]$
$=\prod_{i=1}^{n}\mathbb{P}[\mathcal{R}(P)_{i}=x_{i}]$
$=\prod_{i=1}^{n}p_{i,x_{i}}$ (9)
where the second equation holds due to Definition 3. By substituting Eq. (8) and Eq. (9) into Eq. (A), we can have
$\mathop{\mathbb{E}}_{X\sim\mathcal{S}_{P}}[\mathcal{F}(X)]=\sum_{\alpha\in{}D}\hat{\mathcal{F}}(\alpha)\cdot\prod_{i=1}^{n}p_{i,x_{i}}$
Proof of Theorem 3
  • $\Rightarrow$”: Suppose the Boolean formula $F$ is satisfiable and $X\in D$ is one of the solutions. Then a real assignment $P\in\tilde{D}$ such that $\mathbb{P}(\mathcal{R}(P)=X)=1$ we have $\mathop{\mathbb{E}}_{X\sim\mathcal{S}_{P}}[f_{c}(X)]=1$ for $c\in{C}$. Therefore $\mathcal{F}(X)=|C|$ and $\max\mathop{\mathbb{E}}_{X\sim\mathcal{S}_{P}}[\mathcal{F}(X)]=|C|$.
  • $\Leftarrow$”: Suppose Eq. (2) holds. For every $c\in{}C$ we have $\mathbb{E}_{X\sim{}\mathcal{S}_{P}}[f_{c}(X)]=1$. Thus, the rounding $X\sim{}\mathcal{S}_{P}$ is a discrete assignment that satisfies all constraints.
Proof of Corollary 1
Let $P\in\tilde{D}$ and $X\in{}D$. For each $c\in{}C$, by linearity of expectation, we have
$\mathop{\mathbb{E}}_{X\sim\mathcal{S}_{P}}[f_{c}(X)]$ $=1\cdot\mathop{\mathbb{P}}_{X\sim\mathcal{S}_{P}}[f_{c}(X)=1]+0\cdot\mathop{\mathbb{P}}_{X\sim\mathcal{S}_{P}}[f_{c}(X)=0]$
$=\mathop{\mathbb{P}}_{X\sim\mathcal{S}_{P}}[f_{c}(X)=1]$
$=\texttt{COP}_{c}(P).$
Hence,
$\mathop{\mathbb{E}}_{X\sim\mathcal{S}_{P}}[\mathcal{F}(X)]$ $=\sum_{c\in{}C}\mathop{\mathbb{E}}_{X\sim\mathcal{S}_{P}}[f_{c}(X)]$
$=\sum_{c\in{}C}\texttt{COP}_{c}(P)$
$=\mathcal{C}(P)$
Proof of Lemma 1
We prove by induction on the MDD $G(V,E)$. Let ${P}\in\tilde{D}$.
  • Basic step: For the root node $r$ of $M$, $m_{td}[r]=1$. This equation holds because $r$ is on the path generated by every discrete assignment on $G$.
  • Inductive step: For each non-root node $v$ of $V$, let $par_{i}(v)$ be the set of parent nodes of $v$ with an edge labeled by $i$. After Algorithm 2 terminate, we have
    $m_{td}[v]=\sum_{u\in par_{i}(v)}{p}_{id(u),i}\cdot m_{td}[u],$
    where $id(u)$ is the mapping of the node $u$ to the corresponding variable index. By inductive hypothesis,
    $m_{td}[v]$ $=\sum_{u\in par_{i}(v)}\mathbb{P}[b_{id(u)}=i]\cdot\mathbb{P}[u|G,{P}]$
    $=\mathbb{P}[v|G,{P}]$
    The second equation holds because the events of reaching $v$ from different parents are exclusive.
Proof of Lemma 2
We prove this by structural induction on MDD $G(V,E)$. Let ${P}\in\tilde{D}$, we have
  • Basic step: For the terminal nodes, $m_{bu}[\texttt{true}]=1$ and $m_{bu}[\texttt{false}]=0$. The statement holds since the sub-functions given by terminal nodes are constant functions with values 1 and 0, respectively.
  • Inductive step: For each non-terminal node $v$, let $v.i$ be the child node connecting $v$ by edge label $i$. after algorithm 3 terminates, we have
    $m_{bu}[v]=\sum_{i\in\mathrm{dom}(x_{id(v)})}{p}_{id(v),i}\cdot m_{bu}[v.i].$
    By the inductive hypothesis, we have
    $m_{bu}[v]$ $=\sum_{i=0}^{k}{p}_{id(v),i}\cdot\texttt{COP}_{c,v.i}({P})$
    $=\sum_{i=0}^{k}\mathbb{P}[B_{id(v)}=i]\cdot\texttt{COP}_{c,v.i}({P})$
    $=\texttt{COP}_{c,v}({P})$
    where $\texttt{COP}_{c,v}$ denotes the circuit-output probability of subgraph of $G$ rooted at node $v$.
Proof of Theorem 4
Given an MDD $G(V,E)$, expand all terminal nodes such that each can only have one parent node. The expanded MDD is called $\mathcal{G}$, which has a true terminal set $T$ and a false terminal set $F$. We have the following by repeating the proof of Lemma 1 and 2 on $\mathcal{G}$.
$\texttt{COP}_{c}({P})=\sum_{v\in{}T}\mathbb{P}[v|{P},\mathcal{G}]=\sum_{v\in{}T}m_{td}^{\prime}(v)$
For any non-terminal node $v\notin(T\cup F)$,
$m_{td}[v]$ $=m_{td}^{\prime}[v],$
$m_{bu}[v]$ $=m_{bu}^{\prime}[v].$
For any true terminal node $v\in{T}$ and false terminal node $u\in{F}$,
$m_{bu}^{\prime}[v]$ $=1,$
$m_{bu}^{\prime}[u]$ $=0.$
Now, we are ready to prove Theorem 4.
$\texttt{COP}_{c}({P})$ $=\mathbb{P}[f_{c}=1|{P},\mathcal{G}]$
$=\sum_{v\in\{u:id(u)=i\}}\mathbb{P}[v|{P},\mathcal{G}]\cdot\mathbb{P}[f_{c}=1|v,{P},\mathcal{G}]$
$=\sum_{v\in\{u:id(u)=i\}}m_{td}^{\prime}[v]\cdot m_{bu}^{\prime}[v]$
$=\sum_{v\in\{u:id(u)=i\}}m_{td}[v]\cdot m_{bu}[v]$
$=\sum_{v\in\{u:id(u)=i\}}m_{td}[v]\cdot\sum_{j=0}^{k}p_{i,j}m_{bu}[v.j]$
Hence we have
$\frac{\partial\texttt{COP}_{c}(P)}{\partial p_{i,j}}=\sum_{v\in\{u:id(u)=i\}}m_{td}[v]\cdot m_{bu}[v.j]$
Proof of Lemma 3
To prove Lemma 3, we first establish three propositions that collectively support the main result.
Proposition 1 (Lipschitz).
Let $\mathcal{C}:\tilde{D}\to[0,|C|]$ be a multilinear polynomial. Let $n$ be the number of variable, then $N=\sum_{i=1}^{n}|\mathrm{dom}(x_{i})|$. For every point $y,z\in\tilde{D}$, we have
$|\mathcal{C}(y)-\mathcal{C}(z)|$ $\leq|C|N^{\frac{1}{2}}\|y-z\|,$
i.e., $\mathcal{C}$ is $|C|N^{\frac{1}{2}}$-Lipschitz continuous, and $L=|C|N^{\frac{1}{2}}$ is the Lipshitz constant. As a consequence of the Lipschitz continuity, we have
$|\mathcal{C}(y)-\mathcal{C}(z)-\left<\nabla\mathcal{C}(z),y-z\right>|\leq\frac{L}{2}\|y-z\|^{2}.$ (10)
Proof of Proposition 1.
By the triangle inequality,
$|\mathcal{C}(y)-\mathcal{C}(z)|$
$\leq$ $|\mathcal{C}(y_{1,1},y_{1,2},\cdots,y_{n,|\mathrm{dom}(x_{n})|})-\mathcal{C}(z_{1,1},y_{1,2},\cdots,y_{n,|\mathrm{dom}(x_{n})|})|$
$+$ $|\mathcal{C}(z_{1,1},y_{1,2},\cdots,y_{n,|\mathrm{dom}(x_{n})|})-\mathcal{C}(z_{1,1},z_{1,2},\cdots,y_{n,|\mathrm{dom}(x_{n})|})|+\cdots$
$+$ $|\mathcal{C}(z_{1,1},z_{1,2},\cdots,y_{n,|\mathrm{dom}(x_{n})|})-\mathcal{C}(z_{1,1},z_{1,2},\cdots,z_{n,|\mathrm{dom}(x_{n})|})|$
$=$ $|(y_{1,1}-z_{1,1})\partial_{1,1}\mathcal{C}(y_{1,2},\cdots,y_{n,|\mathrm{dom}(y_{i})|})|+\cdots+|(y_{n,|\mathrm{dom}(x_{n})|}-z_{n,|\mathrm{dom}(x_{n})|})\partial_{n,|\mathrm{dom}(x_{n})|}\mathcal{C}(z_{1,1},\cdots,z_{n,|\mathrm{dom}(x_{n})|-1})|$
Due to Theorem 4, $|\partial_{i,j}\mathcal{C}|\leq|C|$ for all $i\in[n]$ and $j\in[|\mathrm{dom}(x_{i})|]$, we have
$|(y_{1,1}-z_{1,1})\partial_{1,1}\mathcal{C}(y_{1,2},\cdots,y_{n,|\mathrm{dom}(y_{i})|})|+\cdots+|(y_{n,|\mathrm{dom}(x_{n})|}-z_{n,|\mathrm{dom}(x_{n})|})\partial_{n,|\mathrm{dom}(x_{n})|}\mathcal{C}(z_{1,1},\cdots,z_{n,|\mathrm{dom}(x_{n})|-1})|$
$\leq$ $|C|\left(\sum_{i\in[n]}\sum_{j\in[|\mathrm{dom}(x_{i})|]}|y_{i,j}-z_{i,j}|\right)$
$\leq$ $|C|N^{\frac{1}{2}}\|y-z\|$
The last inequality is an application of the Cauchy-Schwarz Inequality.
The proof of Eq. (10) is given by representing $\mathcal{C}(y)-\mathcal{C}(z)$ as an integral.
[…]
$\leq$ $\int_{0}^{1}\|\nabla\mathcal{C}(z+t(y-z))-\nabla\mathcal{C}(z)\|\cdot\|y-z\|dt$
$\leq$ $\int_{0}^{1}L\cdot t\cdot\|y-z\|^{2}dt$
$=$ $\frac{L}{2}\|y-z\|^{2}dt$
The Lipschitz continuity bound the smoothness of $\mathcal{C}$. By performing projected gradient descent, $\mathcal{C}$ will converge to the boundary of the simplices. For each projection step, we have the following Proposition.
Proposition 2 (Projection Point (Nesterov 2013)).
For $y,z\in\tilde{D}$, we define the distance function be
$h(y)=\|y-(z+\eta\nabla\mathcal{C}(z))\|^{2},$
where $\eta>0$ is the step size. Let $y^{*}=\operatorname*{arg\,min}_{y\in\tilde{D}}h(y)$ be the Euclidean projection of the ascent step $z+\eta\nabla\mathcal{C}(z)$ onto $\tilde{D}$. Then the following optimality condition holds:
$\left<\nabla h(y^{*}),y-y^{*}\right>\geq 0.$
In particular, the projection point satisfies $y^{*}=z+\eta{}g(z)$, where $g(z)$ denotes the projected gradient mapping in the ascent direction. Hence, we have the following Proposition.
Proposition 3 (Gradient Mapping).
Given a function $\mathcal{C}$ defined on $\tilde{D}$, for every $y\in\tilde{D}$, we have
$\left<\nabla\mathcal{C}(y),g(y)\right>\geq\|g(y)\|^{2},$
where $g(y)$ is the gradient mapping.
Proof of Proposition 3.
Given any point ${z}\in{\tilde{D}}$, the projection point $y^{*}={z}+\eta{}g({z})$. By substituting $y^{*}$ with ${z}+\eta{}g({z})$ in Proposition 2, for all $y\in\tilde{D}$ we have
$\left<\nabla{}h({z}+\eta{}g({z})),y-({z}+\eta{}g({z}))\right>\geq 0.$
Let y=zy={z}, the above inequality becomes
$\left<\nabla{}h(y+\eta{}g(y)),-\eta{}g(y))\right$> $\geq 0$
$\left<2((y+\eta{}g(y))-(y+\eta{}\nabla\mathcal{C}(y))),-\eta{}g(y)\right$> $\geq 0$
$\left<2\eta{}(g(y)-\nabla\mathcal{C}(y)),-\eta{}g(y)\right$> $\geq 0$
$\left<\eta^{2}(\nabla\mathcal{C}(y)-g(y)),g(y)\right$> $\geq 0.$
For any $\eta>0$, we have $\left<(\nabla\mathcal{C}(y)-g(y)),g(y)\right>\geq 0$
By combining the propositions, we are now ready to quantitatively measure the progress made in every gradient step.
Proof of Lemma 3.
Eq. 10 can be rewritten as
$\mathcal{C}(y)\geq\mathcal{C}(z)+\left<\nabla\mathcal{C}(z),y-z\right>-\frac{L}{2}\|y-z\|^{2}$
Let $y=z+\eta{}g(z)$, the above inequality becomes
$\mathcal{C}(y)\geq\mathcal{C}(z)+\eta\left<\nabla\mathcal{C}(z),g(z)\right>-\frac{\eta^{2}\cdot L}{2}\|g(z)\|^{2}.$
Applying Proposition 3 and let $\eta\leq{}\frac{1}{L}$, we then have
$\mathcal{C}(y)$ $\geq\mathcal{C}(z)+\eta\|g(z)\|^{2}-\frac{\eta}{2}\|g(z)\|^{2}$
$=\mathcal{C}(z)+\frac{\eta}{2}\|g(z)\|^{2}.$
Proof of Theorem 5
From Lemma 3, we can assume $\eta=\frac{1}{L}$. Consider a sequence of points ${P}_{t}\in\tilde{D}$ such that ${P}_{t+1}={P}_{t}+\eta g({P}_{t})$. Due to Lemma 3 we have $\mathcal{C}({P}_{(t+1)})-\mathcal{C}({P}_{t})\geq\frac{\eta}{2}\|g({P}_{t})\|^{2}$. Given any initial point $a_{(0)}$, it will converge to ${P}^{*}$. After $T$ gradient step, we have
$\mathcal{C}({P}^{*})-\mathcal{C}({P}_{0})$ $\geq\frac{1}{2L}\sum_{t=0}^{T}\|g(a_{(t)})\|^{2}$
$\geq\frac{T+1}{2L}\min_{t}\|g(a_{(t)})\|^{2},$
which can be rewritten as
$\min_{t}\|g(a_{(t)})\|$ $\leq\sqrt{\frac{2L}{T+1}}\sqrt{\mathcal{C}({P}^{*})-\mathcal{C}({P}_{0})}$
$\leq\sqrt{\frac{2L}{T+1}}\sqrt{|C|}$
To achieve a point ${P}_{T}$ such that $\|g({P}_{t})\|\leq\epsilon$, we require $\sqrt{\frac{2|C|L}{T+1}}\leq\epsilon$, which implies $T\sim\mathcal{O}\left(\frac{|C|L}{\epsilon^{2}}\right)$.
Proof of Lemma 4 and Theorem 6
Since $\mathcal{C}$ is a multilinear function, it can be decomposed as
$\mathcal{C}(A)=\sum_{j\in\mathrm{dom}(x_{1})}a_{1,j}{}g_{j}(A_{\backslash{}1})+h(A_{\backslash{}1}),$
where $g_{j}$’s and $h$ are multilinear polynomials not depending on $j$. Let
$\eta q^{+}_{i}=\begin{cases}{{proj}_{\tilde{D}_{i}}(a_{1}+\eta\nabla\mathcal{C}(a_{1}))-a_{1}},&i=1,\\ \vec{0},&otherwise,\end{cases}$
and $\epsilon$ be an arbitrary positive number. For all $\eta$ such that $\eta\in(0,\epsilon)$, we have
$\mathcal{C}(A+\eta{}q^{+})-\mathcal{C}(A)$
$=$ $\sum_{j\in\mathrm{dom}(x_{1})}\eta q^{+}_{1,j}\cdot g_{j}(A_{\backslash{}1})$
$=$ $\sum_{j\in\mathrm{dom}(x_{1})}\left({proj}_{\tilde{D}}(a_{1}+\eta\nabla\mathcal{C}(a_{1}))_{j}-a_{1,j}\right)\cdot g_{j}(A_{\backslash{}1}).$
If ${proj}_{\Delta_{1}}(a_{1}+\eta{}q_{1}^{+})$ is on $\Delta\backslash{\delta\Delta}$, we have
$\frac{1}{\eta}\left(\mathcal{C}(A+\eta{}q^{+})-\mathcal{C}(A)\right)=$ $\frac{1}{\eta}\sum_{j\in\mathrm{dom}(x_{1})}\left(\eta g_{j}(A_{\backslash{}1})-\frac{\sum_{j\in\mathrm{dom}(x_{1})}\left(a_{1,j}+\eta g_{j}(a_{\backslash{}1})\right)-1}{|\mathrm{dom}(x_{1})|}\right)g_{j}(A_{\backslash{}1})$
$=$ $\sum_{j\in\mathrm{dom}(x_{1})}\left(g_{j}(A_{\backslash{}1})-\frac{\sum_{j\in\mathrm{dom}(x_{1})}g_{j}(A_{\backslash{}1})}{|\mathrm{dom}(x_{1})|}\right)\cdot g_{j}(A_{\backslash{}1})$
$=$ $\sum_{j\in\mathrm{dom}(x_{1})}g_{j}(A_{\backslash{}1})^{2}-\frac{1}{|\mathrm{dom}(x_{1})|}\left(\sum_{j\in\mathrm{dom}(x_{1})}g_{j}(A_{\backslash{}1})\right)^{2}$
$=$ $\left(\sum_{j\in\mathrm{dom}(x_{1})}g_{j}(A_{\backslash{}1})^{2}\right)\left(\sum_{j\in\mathrm{dom}(x_{1})}\frac{1}{|\mathrm{dom}(x_{1})|}\right)-\left(\sum_{j\in\mathrm{dom}(x_{1})}\frac{g_{j}(A_{\backslash{}1})}{\sqrt{|\mathrm{dom}(x_{1})|}}\right)^{2}$
$\geq$ $0$
The last inequality is the Cauchy-Schwarz inequality, where the equality holds when $\mathcal{C}(a)$ is a constant.
If ${proj}_{\tilde{D}}(a\pm\eta{}q^{\pm})$ is on $\delta\tilde{D}$}, we can always find an $\epsilon^{\prime}\in(0,\epsilon)$ such that for all $\eta^{\prime}\in(0,\epsilon^{\prime})$, ${proj}_{\tilde{D}}(a\pm\eta^{\prime}q^{\pm})$ is on $\tilde{D}\backslash\delta\tilde{D}$}.
Lemma 4 states that, any interior point $a\in\tilde{D}\backslash{}\delta\tilde{D}$ cannot be a local maximum of $\mathcal{C}(a)$. In contrast, all local maxima must lie on the boundary $\delta\tilde{D}$.
Appendix B Benchmarks and Encodings
Task Scheduling Problem
Task scheduling problems can be useful in compilers for multi-core CPUs. Given a task graph $G(V,E)$, where $V\ni{}v$ is a set tasks, $E\ni(u\to{}v)$ is a set of task dependency. Given $T$ clock cycles and $S$ workers, we define variables $t_{v}\in\{0,\cdots,T-1\}$ and $s_{v}\in\{0,\cdots,S-1\}$ to denote the task $v$ is execute at $t_{v}$ by worker $s_{v}$. Then, the scheduling problem can be formulated as
$F$ $=F_{1}\wedge F_{2},$
$F_{1}$ $=\bigwedge_{(u\to{}v)\in{}E}(t_{u}<t_{v}),$
$F_{2}$ $=\bigwedge_{(u\to{}v)\notin{}E}\left((t_{u}\neq{}t_{v})\vee(s_{u}\neq{}s_{v})\right).$
For each $T\in\{32,64,128,256,512\}$, $S\in\{4,8,16\}$, we choose $|V|={T\cdot S}/{2}$ to generate 10 directed graph $G(V,E)$. Every variable $v\in{}V$ has a variable index $idx(v)$. Between any two variable $u,v\in{}V$ and $idx(v)>idx(u)$, we add an edge $u\to{}v$ with probability $5^{idx(u)-idx(v)}$.
Circuit-Output Probability
For any $(u\to{}v)\in{}E$, the corresponding Fourier expansion, i.e., circuit-output probability of the constraint $c\in{}F_{1}$ can be written as
$\mathbb{P}[t_{u}<t_{v}]$ $=1-\mathbb{P}[t_{u}\geq{}t_{v}]$
$=1-\sum_{i=0}^{T-1}\sum_{j=i+1}^{T-1}\mathbb{P}[t_{u}\geq{}i]\cdot\mathbb{P}[t_{v}=i],$
where $\mathbb{P}[t_{u}\geq{}i]$ can be efficiently computed by cumulative sum.
For any $(u\to{}v)\notin{}E$, the corresponding Fourier expansion, i.e., circuit-output probability of the constraint $c\in{}F_{2}$ can be written as
$\mathbb{P}[(t_{u}\neq{}t_{v})\vee(s_{u}\neq{}s_{v})]=$ $1-\mathbb{P}[(t_{u}=t_{v})\wedge(s_{u}=s_{v})]$
$=$ $1-\mathbb{P}[(t_{u}=t_{v})]\cdot\mathbb{P}[(s_{u}=s_{v})].$
Linearization
While encoding $F_{1}$ into ILP is straightforward, encoding $F_{2}$ requires auxiliary variables $q_{u,v,i}\in\{0,1\}$.
$F^{int}$ $=F_{1}^{int}\wedge F_{2}^{int}\wedge F_{3}^{int}\wedge F_{4}^{int}\wedge F_{4}^{int}\wedge F_{6}^{int},$
$F_{1}^{int}$ $=\bigwedge_{(u\to{}v)\in{}E}(t_{v}-t_{u}\geq{}1),$
$F_{2}^{int}$ $=\bigwedge_{(u\to{}v)\notin{}E}(t_{u}-t_{v}+Tq_{u,v,1}\geq 1),$
$F_{3}^{int}$ $=\bigwedge_{(u\to{}v)\notin{}E}(t_{v}-t_{u}+Tq_{u,v,2}\geq 1),$
$F_{4}^{int}$ $=\bigwedge_{(u\to{}v)\notin{}E}(s_{u}-s_{v}+Sq_{u,v,3}\geq 1),$
$F_{5}^{int}$ $=\bigwedge_{(u\to{}v)\notin{}E}(s_{v}-s_{u}+Sq_{u,v,4}\geq 1),$
$F_{6}^{int}$ $=\bigwedge_{(u\to{}v)\notin{}E}\left(\sum_{i=1}^{4}q_{u,v,i}\leq 3\right).$
For $F_{2}^{int}$ to $F_{6}^{int}$, the interpretation is as follows: given $(u\to{}v)\notin{}E$, when $(t_{u}=t_{v})\wedge(s_{u}=s_{v})$, all $q_{u,v,i}$ should be 1 to satisfy the constraints in $F_{2}^{int}$ to $F_{5}^{int}$. However, the constraint in $F_{6}^{int}$ is violated. Only when $(t_{u}\neq{}t_{v})\vee(s_{u}\neq{}s_{v})$, at least one of the $q_{u,v,i}$ can be 0, then the constraint in $F_{6}^{int}$ can be satisfied.
Booleanization
To further Booleanize this problem into Boolean constraints, we use a log-2 encoding.
$t_{v}$ $=\sum_{i=0}^{\log_{2}(T)-1}2^{i}t_{v,i},$ (11)
$s_{v}$ $=\sum_{i=0}^{\log_{2}(S)-1}2^{i}s_{v,i},$ (12)
where $t_{v,i}\in\{0,1\}$ and $s_{v,i}\in\{0,1\}$. Similarly, substituting $F_{1}^{int}$ with Eq. (11) directly leads to $F_{1}^{bool}$. And for encoding $F_{2}$, we introduce $q_{u,v,i}^{(t)}\in\{0,1\}$ and $q_{u,v,i}^{(s)}\in\{0,1\}$.
$F^{bool}$ $=F_{1}^{bool}\wedge F_{2}^{bool}\wedge F_{3}^{bool},$
$F_{1}^{bool}$ $=\bigwedge_{(u\to{}v)\in{}E}\left(\sum_{i=0}^{\log_{2}(T)-1}2^{i}t_{v,i}-\sum_{i=0}^{\log_{2}(T)-1}2^{i}t_{u,i}\geq{}1\right),$
$F_{2}^{bool}$ $=\bigwedge_{(u\to{}v)\notin{}E}\bigwedge_{i=0}^{\log_{2}(T)-1}(t_{u,i}\oplus t_{v,i}\oplus\neg{}q_{u,v,i}^{(t)}),$
$F_{3}^{bool}$ $=\bigwedge_{(u\to{}v)\notin{}E}\bigwedge_{i=0}^{\log_{2}(S)-1}(s_{u,i}\oplus s_{v,i}\oplus\neg{}q_{u,v,i}^{(s)}),$
$F_{4}^{bool}$ $=\bigwedge_{(u\to{}v)\notin{}E}\left(\bigvee_{i=0}^{\log_{2}(T)-1}q_{u,v,i}^{(t)}\vee\bigvee_{i=0}^{\log_{2}(S)-1}q_{u,v,i}^{(s)}\right).$
$F_{2}^{bool}$ and $F_{3}^{bool}$ (resp. $F_{4}^{bool}$) should be interpreted similarly to $F_{2}^{int}$ to $F_{5}^{int}$ (resp. $F_{6}^{int}$). Any $t_{u,i}\neq t_{v,i}$ will lead to $q_{u,v,i}^{(t)}=1$ (or $s$ equivalently), the constraint in $F_{4}^{bool}$ will be satisfied.
Graph Coloring Problem with Hashing Queries
Given a graph $G(V,E)$, where $V\ni{}v$ is a vertex set, $E\ni(u,v)$ is an edge set. Given $C$ colors, we define variables $x_{v}\in\{0,\cdots,C-1\}$ to denote the vertex $v$ is colored by $x_{v}$.
Then, the graph coloring problem can be formulated as a constraint set
$C_{1}$ $=\bigwedge_{(u,v)\in{}E}(x_{u}\neq{}x_{v}).$
In this problem, we additionally use a hash function to generate soft modulo constraints. The hash code $h\in{}H$ randomly samples half of the variables from the vertex set. The hash function is defined by $\sum_{i\in h}x_{i}\mod 2$. Then, the soft constraint set is:
$C_{2}$ $=\bigwedge_{h\in{}H}\left(\sum_{i\in h}x_{i}\mod 2=1\right).$
The constraint optimization problem is
$\min_{a}|H|\cdot\text{cost}(C_{1},a)+\text{cost}(C_{2},a),$
where $\text{cost}(C,a)$ is the number of unsatisfied constraints in $C$ when given an assignment $a$.
For each $|V|\in\{512,1024,2048,4096\}$, $C\in\{8,16,32,64\}$, we generate 10 random regular graphs with degree $2C$. We add $|V|$ modulo constraints as soft constraints.
Circuit-Output Probability
For any $(u,v)\in{}E$, the corresponding Fourier expansion, i.e., circuit-output probability of the constraint $c\in{}C_{1}$ can be written as
$\mathbb{P}[x_{u}\neq{}x_{v}]$ $=1-\mathbb{P}[x_{u}=x_{v}]$
$=1-\mathbb{P}[x_{u}=i]\cdot\mathbb{P}[x_{v}=i].$
For any $h\in{}H$, as inspired by the Fourier expansion of XOR constraints (Kyrillidis et al. 2020), the circuit-output probability of the constraint $c\in{}C_{2}$ can be written as
[…]
$=$ $\frac{1-\prod_{i\in{}h}\left(1-2\sum_{j=0}^{\frac{C}{2}-1}\mathbb{P}[x_{i}=2j+1]\right)}{2},$
which can be handled by slicing.
After the above relaxations, the main computation of FourierCSP can be amenable for parallelization using Jax (Bradbury et al. 2018).
Linearization
When encoding $C_{1}$ for ILP, we introduce two auxiliary variables $q_{u,v},q_{v,u}\in\{0,1\}$ for every $(u,v)\in{}E$.
$C_{1}^{int}$ $=\bigwedge_{(u,v)\in{}E}\left(x_{u}-x_{v}+Cq_{u,v}\geq{}1\right),$
$C_{2}^{int}$ $=\bigwedge_{(u,v)\in{}E}\left(x_{v}-x_{u}+Cq_{v,u}\geq{}1\right),$
$C_{3}^{int}$ $=\bigwedge_{(u,v)\in{}E}\left(q_{u,v}+q_{v,u}\leq{}1\right).$
When encoding $C_{2}$ for ILP, we introduce two auxiliary variables $q_{h}\in\left\{0,...,\frac{|h|(C-1)}{2}\right\}$, and $p_{h}\in\{0,1\}$ for every $h\in{}H$.
$C_{4}^{int}$ $=\bigwedge_{h\in{}H}\left(\sum_{i\in h}x_{i}-2q_{h}+p_{h}=0\right),$
where $p_{h}$ should be 1 (resp. 0) when $\sum_{i\in h}x_{i}$ is odd (resp. even). Then ILP should be formulated as:
$\min|H|-\sum_{h\in{}H}p_{h}$
$\text{s.t. }C_{1}^{int}\cup C_{2}^{int}\cup C_{3}^{int}\cup C_{4}^{int}$
Booleanization
The Booleanization for PB-XOR is similar to the previous section.
$x_{v}$ $=\sum_{i=0}^{\log_{2}(C)-1}2^{i}x_{v,i},$ (13)
where $x_{v,i}\in\{0,1\}$.
For every $(u,v)\in{}E$, and $i\in\{0,...,\log_{2}(C)-1\}$, we introduce an auxiliary variable $q_{u,v,i}\in\{0,1\}$. Consequently, $C_{1}$ can be encoded by
$C_{1}^{bool}$ $=\bigwedge_{(u,v)\in{}E}(x_{u,i}\oplus{}x_{v,i}\oplus{}\neg{}q_{u,v,i}),$
$C_{2}^{bool}$ $=\bigwedge_{(u,v)\in{}E}\left(\bigvee_{i=0}^{\log_{2}(C)-1}q_{u,v,i}\right).$
Note that the parity of $x_{u}$ only depends on $x_{u,0}$. Hence, the soft constraint set $C_{2}$ can be encoded by
$C_{3}^{bool}=\bigwedge_{h\in{}H}\left(\bigoplus_{i\in{}h}x_{i,0}\right)$
Eventually, the constrained optimization problem becomes
$\min_{a}|H|\cdot\left(\text{cost}(C_{1}^{bool},a)+\text{cost}(C_{2}^{bool},a)\right)+\text{cost}(C_{3}^{bool},a).$
Appendix C Convergence Analysis
We selected 36 test cases to analyze convergence behavior. We choose $n\in\{4,8,16\}$ and $m\in\{2,3,4\}$ to generate the test cases, where $n$ is the number of variables and $m$ is the number of states of each variable. The constraint expression consists of the variables and the connectives $\wedge,\vee,\oplus,<,>$, along with parentheses for grouping. Representative examples for $n=4$ and $n=8$ are shown in Fig.3 and Fig.4, respectively.
For all cases, the projected gradient aims to maximize the $m_{td}$ of the true terminal node, or equivalently, minimize $1-m_{td}[true]$. The gradient updates are performed using jaxopt.ProjectedGradient with default parameters. Each run is initialized at the uniform point $\left(\frac{1}{m},\cdots,\frac{1}{m}\right)^{n}$. We record the number of iterations required for convergence and identify whether the final solution lies on the boundary or at a corner of the domain. The results are shown in Table 1,  2, and  3. In every test case, the projected gradient converges to a point $\mathbb{P}[true]=1$. Moreover, each convergent point lies at a corner of the domain, i.e., every variable can be deterministically rounded to a discrete state.
Figure 3: Decision diagrams of the test cases with 4 variables. Left: $c_{1},c_{2},c_{3}$; Mid: $c_{13},c_{14},c_{15}$; Right: $c_{25},c_{26},c_{27}$.
Figure 4: Decision diagrams of the test cases with 8 variables. Left: $c_{5},c_{6},c_{7}$; Mid: $c_{17},c_{18},c_{19}$; Right: $c_{29},c_{30},c_{31}$.
#Variable Name Expression #Iter. On Boundary On Corner
4 $c_{1}$ $x_{1}\land x_{2}\land x_{3}\land x_{4}$ 3 $x_{1},x_{2},x_{3},x_{4}=\{1,1,1,1\}$
$c_{2}$ $x_{1}\lor x_{2}\oplus x_{3}\land x_{4}$ 4 $x_{1},x_{2},x_{3},x_{4}=\{1,1,0,1\}$
$c_{3}$ $x_{1}\land x_{2}\land x_{3}\lor x_{4}$ 2 $x_{4}=\{1\}$
$c_{4}$ $c_{2}\wedge c_{3}$ 4 $x_{1},x_{2},x_{3},x_{4}=\{1,1,0,1\}$
8 $c_{5}$ $x_{1}\land\cdots\land x_{8}$ 6 $x_{1},\ldots,x_{8}=\{1,\ldots,1\}$
$c_{6}$
$x_{1}\oplus x_{2}\land x_{3}\lor x_{4}$
$\land x_{5}\land x_{6}\lor x_{7}\oplus x_{8}$
5 $x_{7},x_{8}=\{1,0\}$
$c_{7}$
$x_{1}\land x_{2}\oplus x_{3}\land x_{4}$
$\lor x_{5}\land x_{6}\oplus x_{7}\land x_{8}$
4 $x_{6},x_{7},x_{8}=\{0,1,1\}$
$c_{8}$ $c_{6}\wedge c_{7}$ 8 $x_{1},\ldots,x_{8}=\{1,1,0,0,1,1,0,1\}$
16 $c_{9}$
$x_{1}\land x_{2}\land\cdots\land x_{16}$
12
$x_{1},\ldots,x_{16}=\{1,\ldots,1\}$
$c_{10}$
$x_{1}\land x_{2}\land x_{3}\lor x_{4}\lor x_{5}\oplus x_{6}$
$\land x_{7}\oplus x_{8}\lor x_{9}\land x_{10}\oplus x_{11}$
$\land x_{12}\lor x_{13}\land x_{14}\lor x_{15}\lor x_{16}$
21 $x_{16}=\{1\}$
$c_{11}$
$x_{1}\oplus x_{2}\land x_{3}\lor x_{4}\land x_{5}\oplus x_{6}$
$\land x_{7}\oplus x_{8}\lor x_{9}\land x_{10}\oplus x_{11}$
$\lor x_{12}\land x_{13}\land x_{14}\oplus x_{15}\land x_{16}$
9 $x_{13},x_{14},x_{15},x_{16}=\{0,0,1,1\}$
$c_{12}$ ${c_{10}}\wedge{c_{11}}$ 7 $x_{13},x_{15},x_{16}=\{0,1,1\}$
Table 1: Convergence analysis of CSP formula where each variable has 2 states.
#Variable Name Expression #Iter On Boundary On Corner
4 $c_{13}$ $x_{1}\land x_{2}\land x_{3}\land x_{4}=1$ 4 $x_{1},x_{2},x_{3},x_{4}=\{1,1,1,1\}$
$c_{14}$ $(x_{1}<x_{2})\oplus(x_{3}>x_{4})=1$ 14 $x_{1},x_{2},x_{3},x_{4}=\{2,0,2,0\}$
$c_{15}$ $x_{1}\land(x_{2}<x_{3})\land x_{4}=1$ 4 $x_{1},x_{2},x_{3},x_{4}=\{1,0,2,1\}$
$c_{16}$ $c_{14}\wedge c_{15}$ 3 $x_{1},x_{2},x_{3},x_{4}=\{1,0,2,1\}$
8 $c_{17}$ $x_{1}\land x_{2}\land\cdots\land x_{8}=1$ 9 $x_{1},\ldots,x_{8}=\{1,\ldots,1\}$
$c_{18}$
$\big(x_{1}\land x_{2}\land x_{3}\lor x_{4}\oplus$
$(x_{5}<x_{6})\land x_{7}\land x_{8}\big)=1$
12 $x_{5},x_{6}$ $x_{4},x_{7},x_{8}=\{1,1,1\}$
$c_{19}$
$\big(x_{1}\lor x_{2}\oplus(x_{3}<x_{4})\lor$
$x_{5}\land(x_{6}<x_{7})\oplus x_{8}\big)=1$
6 $x_{6},x_{7}$ $x_{8}=\{1\}$
$c_{20}$ $c_{18}\wedge c_{19}$ 7 $x_{1},x_{2},x_{3},x_{6}$ $x_{4},x_{5},x_{7},x_{8}=\{1,2,1,1\}$
16 $c_{21}$
$x_{1}\land x_{2}\land\cdots\land x_{16}=1$
108 $x_{1},\ldots,x_{16}=\{1,\ldots,1\}$
$c_{22}$
$\big((x_{1}<x_{2})\land(x_{3}>x_{4})$
$\oplus x_{5}\land(x_{6}<x_{7})\oplus x_{8}$
$\land x_{9}\land x_{10}\lor x_{11}\land x_{12}\oplus$
$x_{13}\lor x_{14}\oplus(x_{15}<x_{16})\big)=1$
139 $x_{15},x_{16}$ $x_{12},x_{13}=\{0,1\}$
$c_{23}$
$\big((x_{1}>x_{2})\land x_{3}\land(x_{4}<x_{5})$
$\lor x_{6}\oplus(x_{7}>x_{8})\land x_{9}$
$\lor x_{10}\lor(x_{11}>x_{12})\lor$
$(x_{13}<x_{14})\land(x_{15}<x_{16})\big)=1$
7 $x_{3},x_{6},x_{9}$ $x_{10},x_{15},x_{16}=\{1,0,2\}$
$c_{24}$ $c_{22}\wedge c_{23}$ 96 $x_{3},x_{6},x_{9}$ $x_{10},x_{12},x_{13},x_{14},x_{15},x_{16}=\{1,0,0,0,0,2\}$
Table 2: Convergence analysis of CSP formula where each variable has 3 states.
#Variable Name Expression #Iter. On Boundary On Corner
4 $c_{25}$ $x_{1}\land x_{2}\land x_{3}\land x_{4}=1$ 5 $x_{1},x_{2},x_{3},x_{4}=\{1,1,1,1\}$
$c_{26}$ $x_{1}\land(x_{2}<x_{3})\land x_{4}=1$ 5 $x_{1},x_{2},x_{3},x_{4}=\{1,1,2,1\}$
$c_{27}$ $x_{1}\land x_{2}\land(x_{3}>x_{4})=1$ 4 $x_{3},x_{4}$ $x_{1},x_{2}=\{1,1\}$
$c_{28}$ $c_{26}\wedge c_{27}$ 3 $x_{3}$ $x_{1},x_{2},x_{4}=\{1,1,1\}$
8 $c_{29}$ $x_{1}\land\cdots\land x_{8}=1$ 11 $x_{1},\ldots,x_{8}=\{1,\ldots,1\}$
$c_{30}$
$\big((x_{1}<x_{2})\land(x_{3}>x_{4})\land$
$x_{5}\lor(x_{6}\oplus x_{7})\land x_{8}\big)=1$
4 $x_{1},x_{2},x_{3},x_{4}$ $x_{5},x_{8}=\{1,1\}$
$c_{31}$
$\big(x_{1}\land x_{2}\oplus x_{3}\lor(x_{4}<x_{5})$
$\oplus x_{6}\land x_{7}\oplus x_{8}\big)=1$
3 $x_{7},x_{8}=\{0,1\}$
$c_{32}$ $c_{30}\wedge c_{31}$ 63 $x_{5}$ $x_{6},x_{7},x_{8}=\{1,0,1\}$
16 $c_{33}$
$x_{1}\land x_{2}\land\cdots\land x_{16}$
145
$x_{1},\ldots,x_{16}=\{1,\ldots,1\}$
$c_{34}$
$x_{1}\oplus(x_{2}>x_{3})\land x_{4}\lor x_{5}$
$\oplus(x_{6}>x_{7})\lor(x_{8}<x_{9})$
$\land x_{10}\land(x_{11}<x_{12})\oplus x_{13}$
$\lor(x_{14}>x_{15})\oplus x_{16}$
52 $x_{11},x_{12}$ $x_{13},x_{16}=\{1,0\}$
$c_{35}$
$\big((x_{1}<x_{2})\land(x_{3}<x_{4})$
$\land x_{5}\oplus x_{6}\land x_{7}\oplus(x_{8}>x_{9})$
$\land x_{10}\land x_{11}\lor x_{12}\lor$
$(x_{13}>x_{14})\land(x_{15}>x_{16})\big)=1$
31 $x_{11},x_{15},x_{16}$ $x_{10},x_{12}=\{0,1\}$
$c_{36}$ $c_{35}\wedge c_{36}$ 41 $x_{15}$ $x_{10},x_{11},x_{12},x_{13},x_{16}=\{0,0,1,1,0\}$
Table 3: Convergence analysis of CSP formula where each variable has 4 states.
Appendix D Parallel Decision Diagram Traversal
The decision diagram can be traversed in parallel. For example, MDD $c_{13}$ and $c_{15}$ can be represented as Listing 1 and Listing 2.
1# vid, nid, eid, cid
21 1 0 5
31 1 1 2
41 1 2 5
52 2 0 5
62 2 1 3
72 2 2 5
83 3 0 5
93 3 1 4
103 3 2 5
114 4 0 5
124 4 1 6
134 4 2 5
140 0 0 0
150 0 0 0
160 0 0 0
Listing 1: c13.mdd
1# vid, nid, eid, cid
21 1 0 6
31 1 1 2
41 1 2 6
52 2 0 3
62 2 1 4
72 2 2 6
83 3 0 6
93 3 1 5
103 3 2 5
113 4 0 6
123 4 1 6
133 4 2 5
144 5 0 6
154 5 1 7
164 5 2 6
Listing 2: c15.mdd
For every MDD, we call the kernel functions to traverse one edge at a time. The corresponding top-down and bottom-up message propagation kernels are shown in Listing 3 and Listing 4. To ensure uniform execution across threads, shorter MDD traces are padded with zeros. The kernel functions are designed to detect such padded entries (e.g., when nid equals zero) and skip message propagation accordingly.
1__global__ void top_down_kernel(
2 const int* variable_id, const int* node_id, const int* edge_id, const int* child_id,
3 const float* p, const int* pointer_id, float* m_td
4) {
5 int idx = blockIdx.x * blockDim.x + threadIdx.x;
6 if (idx >= num_mdd) return;
7
8 int vid = variable_id[idx];
9 if (vid == 0) return;
10 int nid = node_id[idx];
11 int eid = edge_id[idx];
12 int cid = child_id[idx];
13 int pid = pointer_id[vid];
14
15 atomicAdd(&m_td[cid], m_td[nid] * p[pid + eid]);
16}
Listing 3: CUDA-style top-down message propagation kernel function
1__global__ void bottom_up_kernel(
2 const int* variable_id, const int* node_id, const int* edge_id, const int* child_id,
3 const float* p, const int* pointer_id, const int num_mdd, float* m_bu
4) {
5 int idx = blockIdx.x * blockDim.x + threadIdx.x;
6 if (idx >= num_mdd) return;
7
8 int vid = variable_id[idx];
9 if (vid == 0) return;
10 int nid = node_id[idx];
11 int eid = edge_id[idx];
12 int cid = child_id[idx];
13 int pid = pointer_id[vid];
14
15 atomicAdd(&m_bu[nid], m_bu[cid] * p[pid + eid]);
16}
Listing 4: CUDA-style bottom-up message propagation kernel function